# Dynamic Learning Rate for Deep Reinforcement Learning: A Bandit Approach

### v2

@@ B.1 Baseline Evaluation with Varying Learning Rates
To establish a baseline to compare learning without our approach LRRL, we run individual arm values as baseline learning rate using the Adam optimizer. The results presented in Figure 7 align with common expectations by showing that higher learning rates fail to learn for most environments while lower ones can lead to worse jumpstart performance and slow convergence.
Figure 7: DQN performance using the Adam optimizer with varying learning rates across 6 Atari games. Shaded region represents one-half standard deviation. The results show that the best performing learning rate varies across games.
B.2 A Comparison between Multi-Armed Bandits Algorithms
Adversarial MAB algorithms are designed for environments where the reward distribution changes over time. In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown probability distributions. To validate our choice and address our method’s robustness, we compare Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [5]. MOSS trades off exploration and exploitation by pulling the arm $k$ with highest upper confidence bound given by:
@@ B.2 A Comparison between Multi-Armed Bandits Algorithms
where $\hat{\mu}_{k}(n)$ is the empirical average reward for arm $k$ and $n_{k}(n)$ is the number of times it has been pulled up to round $n$.
In Figure 8, we use different bandit step-sizes $\alpha$ for Exp3 and parameter $\rho$, which balances exploration and exploitation in MOSS. Additionally, the bandit feedback used in MOSS is the average cumulative reward $\frac{R}{C}$. The results indicate that while Exp3 performs better overall, MOSS can still achieve competitive results depending on the amount of exploration, demonstrating the robustness of our approach regarding the MAB algorithm employed.
Figure 8: A comparison between adversarial (Exp3) and stochastic (MOSS) MAB algorithms. Shaded region represents one-half standard deviation. The results show that Exp3 performs better overall.
B.3 Dealing with Noise and Lack of Feedback
LRRL relies entirely on performance improvements to update the arm probabilities. However, some RL environments can be highly stochastic, with sparse or delayed rewards, which increases the difficulty of learning tasks and, consequently, the ability of LRRL to make optimal choices. To mitigate this limitation, we introduce the variable $\kappa$ in Algorithm 1, which determines the number of interactions to observe before making a decision.
When setting $\kappa$, two factors should be considered:
  • The network’s update rate: Selecting new learning rates more frequently than updating the model leads to inaccurate feedback for the bandit mechanism since no model update might occur after pulling some arm.
  • The number of environment steps: In some tasks, positive or negative outcomes of a sequence of actions are only perceived at the end of an episode, and capturing sufficient information about the agent performance is essential to updating the arms’ probabilities.
In the experiments presented in this work, we set $\kappa=1$ episode, allowing the observation of a complete episode before updating the arms’ probabilities. To explore the impact of this parameter further, we conduct an ablation study using different $\kappa$ values to evaluate potential performance improvements when combined with the schedulers shown in Figure 3. As suggested by Figure 9, the choice of $\kappa$ is task-dependent: while some tasks benefit from more reactive learning rate updates, others show marginal improvements with increased observation counts. Ultimately, setting $\kappa=1$ seems relatively robust for Atari games, although higher values could slightly increase the performance in some tasks.
B.4 Supplementary Results using SGD
In Section 4.5, we tested the SGD optimizer in $6$ non-convex loss landscapes show in Figure 10. These functions were selected since they represent common benchmarks in optimization, each one offering different challenges due to their distinct shapes. By using SGD on these landscapes, we minimize the influence of confounding factors, such as the high stochasticity inherent in some RL tasks and the complexity of more robust optimizers like Adam.
Figure 11 presents the normalized loss and the arms selected at each optimization step. The results demonstrate LRRL’s dynamic adaptation mechanism, either converging to the most effective learning rate or combining them to achieve superior outcomes.
@@ B.4 Supplementary Results using SGD
Figure 10: The 6 stationary non-convex loss landscapes used to assess LRRL with SGD optimizer.
Figure 11: Learning rates and logarithmic scale of losses over the gradient steps using LRRL with SGD optimizer.
Appendix C Evaluation Metrics
@@
In this section, we describe the evaluation metrics that can be used to evaluate the agent’s performance as it interacts with an environment. Based on Dopamine, we use the evaluation step-size “iterations”, defined as a predetermined number of episodes. Figure 12 illustrates the evaluation metrics used in this work, as defined in Taylor and Stone [40]:
@@
Figure 12: Performance curves of two RL algorithms (adapted from 40).
Appendix D Implementation Details
@@
Section 4.1 – Comparing LRRL with Standard Learning.
@@
    $\mathcal{K}(5)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
    $\mathcal{K}_{\text{lowest}}(3)$ =$=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
    $\mathcal{K}_{\text{middle}}(3)$ =$=\hskip 65.44142pt\bigl{\{}3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr{\}}$
    $\mathcal{K}_{\text{highest}}(3)$ =$=\hskip 125.19194pt\bigl{\{}6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr{\}}$
@@
Section 4.2 – Combining and Comparing Schedulers with LRRL.
@@
Section 4.3– Comparing RMSProp and Adam Optimizers
@@
  • LRR-DQN set of learning rates: $\mathcal{K}=\bigl{\{}1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr{\}}$
Section 4.4– LRRL with a Distributional RL Algorithm.
@@
Section B.2 – A Comparison between Multi-Armed Bandit Algorithms.
  • Optimizer: Adam
@@
For reproductibility, the list of hyperparameters of the optimizers and deep RL algorithms used for our experiments is provided in Table 1. The optimizers from the Optax library [12] are employed alongside the JAX [8] implementation of the DQN [25, 26] and IQN [11] algorithms, as provided by the Dopamine framework [9].
@@
Eps 1.5$\times$10-4
Table 1: Hyperparameters used in the experiments.

### v3

@@ B.1 Baseline Evaluation with Varying Learning Rates
To establish baselines, we evaluate DQN with Adam across several Atari tasks using fixed learning rates. As shown in Figure 7, larger learning rates may provide a better jumpstart but often lead to worse asymptotic performance, while smaller rates result in more stable convergence but slower learning, with the best-performing learning rate varying by environment. This illustrates why a dynamic strategy such as LRRL, which is able to exploit higher rates early and smaller ones later, can be advantageous.
Figure 7: Baseline evaluation of DQN with Adam using different fixed learning rates across Atari tasks. Performance varies substantially across games, confirming that no single learning rate works universally.
B.2 Comparing adversarial and stochastic MAB algorithms
Adversarial MAB algorithms are designed for settings where reward distributions can change over time, making them naturally suited to RL. In contrast, stochastic MAB algorithms assume that rewards are drawn from fixed but unknown distributions. To validate our design choice and assess robustness, we compare LRRL using Exp3 with the stochastic MAB algorithm MOSS (Minimax Optimal Strategy in the Stochastic case) [4].
MOSS balances exploration and exploitation by selecting the arm $k$ with the largest upper confidence bound:
@@ B.2 Comparing adversarial and stochastic MAB algorithms
where $\hat{\mu}_{k}(n)$ is the empirical average reward of arm $k$, and $n_{k}(n)$ is the number of times it has been pulled up to round $n$. In our experiments (Figure 8), we vary the step-size $\alpha$ for Exp3 and the exploration parameter $\rho$ for MOSS, using average cumulative return $f_{n}$ as feedback for MOSS.
Results show that Exp3 achieves stronger overall performance, consistent with the non-stationary nature of RL. Nevertheless, MOSS can reach competitive outcomes under appropriate exploration settings, demonstrating that LRRL is robust to the choice of bandit algorithm.
Figure 8: Comparison of adversarial (Exp3) and stochastic (MOSS) bandit algorithms within LRRL. Exp3 performs better overall, as expected in non-stationary RL settings, but MOSS can achieve competitive results under certain exploration parameters, highlighting robustness to the bandit choice.
B.3 Comparing RMSProp and Adam Optimizers
We next evaluate LRRL’s robustness to the choice of optimizer. We compare Adam and RMSProp augmented with Nesterov momentum (RMSProp-M), since prior work [3] and our own observations indicate that adding momentum consistently improves RMSProp’s performance. Both Adam and RMSProp-M are paired with either LRRL and the best-performing fixed learning rate within the DQN agent. Since LRRL has already shown consistency in earlier sections, we extend the evaluation to three additional Atari games here.
As shown in Figure 9, LRRL combined with Adam achieves the strongest results overall, outperforming both RMSProp-M baselines and the tuned fixed-rate Adam baseline. LRRL with RMSProp-M provides better jumpstart but converges more slowly, underperforming standard DQN in two out of the three tasks. One possible explanation is the lack of bias correction in RMSProp’s moment estimates, which may limit LRRL’s ability to fully adapt. Exploring such optimizer-specific interactions is an interesting direction for future work.
Figure 9: Comparison of Adam and RMSProp with Nesterov momentum (RMSProp-M), each combined with either LRRL or the best-performing fixed learning rate on DQN across three additional Atari games. LRRL with Adam consistently achieves the strongest results, while LRRL with RMSProp-M shows good jumpstart performance but slower convergence.
B.4 Distributional RL: Implicit Quantile Networks (IQN)
To further illustrate the generality of LRRL, we combine it with Implicit Quantile Networks (IQN) [12], a distributional RL algorithm available in Dopamine that has shown stronger empirical performance than DQN and other algorithms in several Atari tasks. Figure 10 shows that LRRL improves or matches performance across the evaluated Atari tasks when paired with IQN. These results demonstrate that LRRL is not tied to a specific algorithmic family: beyond standard value-based methods such as DQN or actor–critic methods such as PPO, LRRL can also enhance distributional agents. More broadly, since LRRL operates at the optimizer level by dynamically selecting the learning rate, it can in principle be coupled with any algorithm based on SGD or its variants. This makes LRRL widely applicable across deep RL.
Figure 10: IQN with Adam on Atari tasks: LRRL vs. fixed learning rates. LRRL improves or matches performance across environments, showing that dynamic learning-rate adaptation benefits stronger distributional RL agents as well as DQN.
B.5 Supplementary Results using SGD
@@
Figure 11: The six stationary non-convex loss landscapes used to assess LRRL with the SGD optimizer. Each function presents distinct challenges, such as narrow valleys (Rosenbrock) or multiple local minima (Griewank).
(a) Beale
(b) Rosenbrock
(c) Three-Hump Camel
(d) Zakharov
Figure 12: Convergence steps toward the global minimum using LRRL and CLR on four benchmark functions.
In Section 4.4, we tested LRRL with SGD on six non-convex loss landscapes (Figure 11). These benchmarks are widely used in optimization because their diverse shapes highlight different challenges, such as multiple local minima, steep ridges, or flat plateaus. Using SGD in this stationary setting minimizes the influence of confounders such as environment stochasticity in RL or the adaptivity of more sophisticated optimizers like Adam.
Here, we extend our experiments and compare LRRL against Cyclical Learning Rates (CLR) [40]. CLR is motivated by a similar intuition that varying learning rates can improve performance, but it follows a predetermined periodic schedule that cycles through candidate values with a fixed step size $n$ (see Algorithm 2). Unlike LRRL, CLR does not use feedback to guide its choices. As a result, while CLR can occasionally reach strong solutions, it may introduce instability by switching rates at arbitrary times. In contrast, LRRL adapts its selection directly from performance feedback.
Algorithm 2 Cyclical Learning Rates [40]
1:Input:
2:$n$: step size (half-cycle length in iterations)
3:minlr: minimum learning rate
4:maxlr: maximum learning rate
5:$M$: total number of training steps
6:for $m=1,2,\dots,M$ do
7:  $\text{cycle}\leftarrow\left\lfloor 1+\frac{m}{2n}\right\rfloor$   (Compute current half-cycle index)
8:  $x\leftarrow\left|\frac{m}{n}-2\cdot\text{cycle}+1\right|$   (Normalized position within cycle)
9:  $\eta\leftarrow\text{minlr}+(\text{maxlr}-\text{minlr})\cdot\max(0,1-x)$   (Compute learning rate)
10:end for
Figure 12 shows the gradient steps taken employing the best performing candidate, in terms of convergence rate, varying respectively the exploration trade-off given by $\rho$ (LRRL) and the step granularity $n$ (CLR). Table 1 displays the convergence steps to global minima considering loss landscapes where that could be achieved with an Euclidean distance less than $0.01$. Notably, LRRL has superior outcomes despite its higher sensitivity to the hyperparameter choice.
Loss Landscape CLR ($n=2$) CLR ($n=5$) CLR ($n=10$) LRRL ($\rho=1$) LRRL ($\rho=2$) LRRL ($\rho=3$)
Beale - 1268 1248 2083 905 907
Rosenbrock 96267 96267 96275 65318 118167 66257
Three-Hump Camel 292 320 359 286 518 169
Zakharov 237 237 236 171 284 178
Table 1: Convergence steps to global minima (Euclidean distance <0.01<0.01) for different loss landscapes and learning rate schedules.
Appendix C Additional Discussion and Limitations
C.1 On Selecting the Learning Rates and Hyperparameters
Learning rates.
Figure 7 shows that different learning rates lead to very different policy performance, yet there is no principled way to select them without extensive testing. LRRL does not eliminate the need to define a candidate set, and its performance naturally depends on this choice, as seen in Figure 1. However, LRRL reduces the burden of tuning by adaptively selecting among candidates within a single run, instead of requiring multiple runs to evaluate each rate individually. As illustrated in the SGD experiments (Figure 6), this allows LRRL to combine aggressive and conservative rates dynamically or to converge toward the best rate when one dominates.
Hyperparameters.
For the bandit algorithm, we fixed $\delta=0.9$ in all experiments, a common choice for decay factors (e.g., momentum or the RL discount factor $\gamma$). For the bandit feedback $f^{\prime}_{n}$, a history window of $j=5$ was robust across environments. As shown in Figure 4, $\kappa$ only slightly affects performance, while $\alpha$ is the most sensitive parameter (Figure 8), requiring tuning. This reflects a broader challenge in meta-optimization: hyperparameters are themselves required to control the process of tuning other hyperparameters [27].
C.2 The Cumulative Reward as a Proxy of Performance
Figure 13: Return and loss in RL.
A central design choice in LRRL is the feedback used by the bandit algorithm. Our setting corresponds to a dual objective: the bandit aims to minimize regret in selecting learning rates, while the RL agent seeks to maximize cumulative returns. We therefore use the improvement in policy performance $f^{\prime}_{n}$ as feedback to update arm probabilities.
Unlike supervised learning, where the training loss directly reflects predictive quality, in RL the optimization loss is only loosely related to policy performance (Figure 13). For example, the TD-error in Equation 1 measures how surprising a transition is, but a high TD-error does not necessarily indicate that the policy is improving. By contrast, returns directly measure policy effectiveness, even if they are noisy or sparse. This noise is intrinsic to RL and affects both LRRL’s feedback and policy optimization itself.
C.3 Computational Overhead
The overhead introduced by LRRL is governed primarily by the decision frequency $\kappa$ (how often we update arm probabilities) and the number of arms $K$. Each bandit update costs $O(K)$ for computing/sampling the distribution, so the total cost over $T$ decisions is $O(TK)$. In our SGD experiments, we update once per gradient step ($\kappa=1$ step), so the extra $O(K)$ work can be noticeable when the model is tiny (e.g., 2D landscapes) because the gradient step itself is very cheap. In RL, we set $\kappa=1$ episode while policy updates happen every $\lambda=4$ steps with mini-batches; thus $T$ is roughly the number of episodes. With small $K$ (as used here), the $O(TK)$ bandit cost is negligible compared to forward/backward passes and replay/batching, and adds no meaningful wall-clock overhead.
C.4 Guarantees
While adversarial bandit algorithms offer theoretical regret guarantees under certain conditions, these guarantees do not carry over to deep RL because the best-performing arm itself changes as the policy evolves. From a theoretical point of view, guarantees would require assumptions (e.g, convexity, smoothness, bounded losses) that deviate significantly from the realities of RL. We therefore focus on empirical validation across diverse benchmarks, demonstrating LRRL’s robustness and generality in practice.
C.5 Future Work
A key limitation of LRRL is the need to predefine a set of candidate learning rates, although our experiments show some robustness to this choice. One possible direction to overcome this limitation is using Bayesian optimization, for example with Gaussian Processes (GPs) [34]. While GPs are sample-efficient, their computational cost increases as the number of observations grows, and they can struggle with non-stationary objectives. Addressing this would likely require mechanisms that emphasize recent interactions, as in time-varying GP-UCB [7], or approaches that limit and/or adapt the number of inducing points [33]. Integrating such techniques with LRRL could reduce reliance on a fixed candidate set while retaining adaptability to non-stationary dynamics.
Appendix D Evaluation Metrics
@@
In this section, we describe the evaluation metrics that can be used to evaluate the agent’s performance as it interacts with an environment. Based on Dopamine, we use the evaluation step-size “iterations”, defined as a predetermined number of episodes. In MuJoCo tasks, updates refer to policy updates. Figure 14 illustrates the evaluation metrics used in this work, as defined in Taylor and Stone [44]:
@@
Figure 14: Performance curves of two RL algorithms (adapted from 44).
Appendix E Implementation Details
@@
Section 4.1 – LRRL vs. Fixed Learning Rates: Robustness Across Candidate Sets
@@
    $\mathcal{K}(5)$ $=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr\}$
    $\mathcal{K}_{\text{lowest}}(3)$ $=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr\}$
    $\mathcal{K}_{\text{middle}}(3)$ $=\hskip 65.44142pt\bigl\{3.125\times 10^{-5},6.25\times 10^{-5},1.25\times 10^{-4}\bigr\}$
    $\mathcal{K}_{\text{highest}}(3)$ $=\hskip 125.19194pt\bigl\{6.25\times 10^{-5},1.25\times 10^{-4},2.5\times 10^{-4}\bigr\}$
@@
Section 4.2 – LRRL with Learning Rate Schedulers.
@@
Section 4.3 – Continuous Action Spaces.
  • Optimizer: Adam
  • Bandit step-size: $\alpha=0.2$
  • AdamRel/Initial learning rate: $3\times 10^{-4}$
  • LRRL/CLR set of learning rates: $\mathcal{K}=\bigl\{1.2\times 10^{-4},1.8\times 10^{-4},2.5\times 10^{-4}\bigr\}$
Section B.2 – Comparing adversarial and stochastic MAB algorithms.
  • Optimizer: Adam
Section B.3– Comparing RMSProp and Adam Optimizers.
@@
  • LRRL-DQN set of learning rates: $\mathcal{K}=\bigl\{1.5625\times 10^{-5},3.125\times 10^{-5},6.25\times 10^{-5}\bigr\}$
Section B.4– Distributional RL: Implicit Quantile Networks (IQN).
@@
For reproductibility, the list of hyperparameters of the optimizers and deep RL algorithms used for our experiments is provided in Table 2. The optimizers from the Optax library [13] are employed alongside the JAX [9] implementation of the DQN [28, 29] and IQN [12] algorithms, as provided by the Dopamine framework [10]. For the MuJoCo tasks, we use the implementation provided by TorchRL [8].
@@
[…]
Table 2: Hyperparameters used in the experiments with discrete action spaces.
PPO hyperparameters Value
Discount factor ($\gamma$) 0.99
Mini-batch size ($\mathcal{B}$) 64
Epochs 10
GAE 0.95
Clip $\epsilon$ 0.2
Critic coef. 0.25
Entropy coef. 0.0
Minimum environment steps ($\kappa$) 1 episode
Network neurons per layer 64, 64
Hardware (GPU) V100
Table 3: Hyperparameters used in the experiments with continuous action spaces.